# A Langevin sampling algorithm inspired by the Adam optimizer

### v1

@@ 5.2.3 Beale Potential
Figure 10: Absolute errors of the means of four observables on the Beale potential against (mean) stepsize $\Delta t$. The ground truths for the coordinates were obtained via numerical quadrature. Hyperparameters: $\gamma=1$, $T=3$, $m=0.1$, $M=10$, $r=0.25$, $\alpha=1$, $g(x,p)=0.1\|\nabla U\|^{2}$. See text for more information.
@@ 5.3 Higher-dimensional Models
We now explore several higher-dimensional cases using fixed stepsize Langevin as well as the SamAdams scheme. Our goal here is to show, using model systems, that there is a good basis for believing the method will be effective for applications in statistics and machine learning. SamAdams has several parameters ($\alpha$, $m$, $M$, $r$, and indeed the monitor function $g$) that can be tuned to adjust the performance of the method. In practice we have found that some attention needs to be paid to their selection to obtain optimal results.
@@
Figure 11: Comparison of trajectory graphs in the $\theta,x_{1}$ projection obtained for different methods and stepsizes. Top row: fixed stepsize BAOAB runs with $\Delta t$ increasing left to right until explosion is encountered at $\Delta t=0.1$. The lower row shows the results of different variable stepsize runs using SamAdams. All of these solutions are usable for sampling purposes, although the last one shows some thickening in the $x$ direction.
We examined the stability of fixed stepsize (BAOAB) Langevin dynamics by running trajectories with a series of increasing timesteps $\Delta t=0.04,0.06,0.08,0.1$. The system was initialised with a large enough value of $\theta$ ($\theta_{0}=5$) and all $x_{i}=0$ so as to avoid any initial high gradient. The graphs in the $\theta,x_{1}$ projection are shown in the upper row of Fig. 11. All simulations involved $N=10^{7}$ steps. We then ran SamAdams with various $\Delta\tau$, generating the figures shown in the second row, also using $10^{7}$ steps for each. Parameters of SamAdams simulations: $m=0.01$, $M=1.0$, $\alpha=1$, and $r=1$, with monitor function
$g(x,p)\equiv\frac{1}{100}\|\nabla U(x)\|$
and filter function $\psi=\psi^{(1)}$. Here $m,M$ define a range of 100 in stepsize, smallest to largest. We label these runs by the mean stepsize used, which is reciprocally related to the computational work to reach a given fixed time. As we can see, all the SamAdams simulations are stable and accurate representations of the true SDE sampling trajectory. Given the mild confining potential, the excursions to large positive $x$ in the fixed stepsize trajectories at stepsize greater than $0.04$ represent high energy (far from equilibrium) events; in a more complicated model those trajectories would lead to blow-up, or, perhaps more seriously, subtle degradation of cumulative averages; here the confining potential quickly returns them to the domain of interest. By contrast, SamAdams is stable and retains the trajectories within a similar size and shape region of the $\theta,x_{1}$ plane, except for a slight increase in the width of the sampled region at the largest mean stepsize ($\langle\Delta t\rangle=0.1762$).
One way to read the results of Fig. 11 is that the largest stable mean stepsize of SamAdams is approximately four times the largest stable stepsize for Langevin dynamics with fixed stepsize.
@@
An important question is whether the time-series corresponding to different timesteps explore the space with similar efficiency. We settle this for the 9D funnel example in Fig. 14 where we see that relative to the elapsed time, the trajectories diffuse at similar rates. [^4] Computing autocorrelation functions and ESS requires first interpolating the non-uniformly spaced time-series data to a uniform mesh; this interpolation can be avoided when computing expectations by using the method of Section 2.1. Effective sample sizes per sample [18] are as shown in Table 13. These results indicate that the trajectories are similar in terms of the rates of exploration.
Mean log posterior values (from 10M step runs) for the different stepsizes are also given in the table and suggest that bias may be becoming more noticeable at the larger stepsize. The result $-10.46$ for mean log posterior is accurate, verified using a small stepsize of $0.01$ and 1 billion steps. Note that with small stepsize $0.01$, fixed stepsize runs of 10M steps as in the table generate similar size errors (in this case due to Monte Carlo error due to the higher correlation of samples) to the largest variable stepsize integration (in that case, due to sampling bias), thus demonstrating a clear trade-off between sampling error and bias. Accuracy (and ESS) appear to fall off at the largest fixed stepsize with $\Delta t=0.02$ being approximately optimal. This can be compared to SamAdams with a mean stepsize of $\langle\Delta t\rangle=0.0998$, indicating an improvement of over 400% in sampling efficiency without any reduction in accuracy.
fixed $\Delta t$ variable $\Delta t$
@@
Figure 13: Table showing the effective sample size per sample and the expected log posterior, for different variable stepsizes.
Figure 14: Autocorrelation functions with respect to iteration number generated by different mean stepsizes. Left: autocorrelation of $\theta$; right: autocorrelation of $x_{1}$. For each variable, the decay rates corresponding to different schemes are similar, regardless of the stepsize (and are similar for both the fixed and variable stepsize runs).
5.3.2 Auxiliary Variable Control Dynamics
Until now we have not yet explored the role of the hyperparameter $\alpha$ in (24) nor the influence of the scale of the monitor function $g(x,p)$. In the case of $g(x,p)=\Omega^{-1}\|\nabla U(x)\|^{r}$ (our choice for all of our experiments in this article), the scaling factor $\Omega$ governs how strongly $\zeta$ reacts to a given force value $\nabla U(x)$. Since the range of typical values of $\nabla U$ strongly depends on the problem at hand, the scaling parameter $\Omega$ is useful to adjust the range of $g(x,p)$ and hence the range of $\zeta$ and $\Delta t$. For the one- and two-dimensional examples considered thus far, setting $\Omega=1$ serves as a good starting point, whereas on the classification tasks of the the next section we use $\Omega=N_{\text{D}}$ with $N_{\text{D}}$ the size of the dataset. Since the force $\nabla U$ is written as a sum over the model likelihoods of all data points, this choice for $\Omega$ encourages $\zeta$-values close to the fixed point of the Sundman transform $\psi$. The hyperparameter $\alpha$ governs the damping of the $\zeta$-dynamics as well as its moving average behavior, i.e., how quickly past force values are ’forgotten’ (see Appendix C for more details).
As we have shown thus far, one fundamental benefit of the adaptive stepsize is the ability to use overall large stepsizes while using small stepsizes selectively. If the resulting $\langle\Delta t\rangle$ is larger than the maximum stable stepsize of a conventional scheme (assuming the accuracies remain acceptable), SamAdams leads to an improvement of stability and computational efficiency. An important question is then how difficult it is to tune the SamAdams-specific hyperparameters $\alpha$ and $\Omega$. To examine this, we chose the example of the star potential. We create a grid of different $(\alpha,\Omega)$-values and run SamAdams trajectories on for each grid point. For all ($\alpha,\Omega$), we experimentally find the stability threshold of SamAdams, i.e., the largest $\langle\Delta t\rangle$ that still leads to a stable trajectory. We do this by running 100 independent trajectories for successively increasing values of $\Delta\tau$, the base stepsize, until the trajectory becomes unstable. We vary $\Delta\tau$ in $[0.04,0.08]$, a range in which the resulting $\langle\Delta t\rangle$-values become roughly comparable to the stability threshold of BAOAB for most ($\alpha,\Omega$)-combinations  [^5] BAOAB’s threshold was found to be $\Delta t=0.01275$, defined as the smallest stepsize for which 100 independent trajectories remain stable for roughly $5\cdot 10^{6}$ iterations.. The number of iterations per tested $\Delta\tau$ is scaled via $N=N_{0}\Delta\tau_{0}/\Delta\tau$ with $N_{0}=2\cdot 10^{6}$ and $\Delta\tau_{0}=0.06$. If an $(\alpha,\Omega)$-combination leads to unstable trajectories even at the smallest $\Delta\tau$ tested, we deem the combination ’highly unstable’. The results are shown in Fig. 15.
Figure 15: SamAdams stability thresholds $\langle\Delta t\rangle_{\text{max}}$ in dependency of $\alpha$ (attack rate) and $\Omega$ (scale coefficient of force norms) in the case of the star potential. Left: Plain $\langle\Delta t\rangle_{\text{max}}$. A well defined optimal zone appears around $\alpha=1$, $\Omega=100$. Right: Fraction of SamAdams threshold and BAOAB’s threshold $\Delta t^{\text{BAOAB}}_{\text{max}}$. The brown areas show parameter regions where BAOAB is more stable than SamAdams, the purple area regions where SamAdams is more stable. The red areas denote ’highly unstable’ regions (see main text). Other hyperparameters: $T=\gamma=1$, $m=0.1$, $M=10$, $r=0.25$, $s=2$, transform kernel $\psi^{(1)}$.
From the left-hand figure, we see how SamAdams’ stability threshold varies with $(\alpha,\Omega)$, with an optimum region visible for $\alpha\simeq 1$ and $\Omega^{-1}\simeq 0.01$. The highly-unstable area (red) is mainly due to too large values of $\alpha$. As shown in Appendix C, too large $\alpha$ will lead to $\Delta t=M\Delta\tau$. For our choice $M=10$ and the smallest $\Delta\tau$-value tested, $\Delta\tau=0.04$, this would correspond to a constant-stepsize scheme running at stepsize 0.4, more than 30 times larger than BAOAB’s stability threshold. Note that even for these highly unstable ($\alpha,\Omega)$-values, one could still enforce stability by picking $\Delta\tau$ and $M$ such that $M\Delta\tau$ is smaller than BAOAB’s threshold. On the right-hand side in Fig. 15, we plot the same grid but color according to the fraction of SamAdams stability threshold and BAOAB’s threshold. We observe that there is a large area in parameter space (shaded purple) in which SamAdams is more stable than BAOAB, with a remarkable improvement of up to 300%.
While these experiments show the need to pick admissible $(\alpha,\Omega)$-values, they also demonstrate that there is a wide range of settings in which SamAdams is more stable than BAOAB, reducing the need for hyperparameter fine-tuning.
@@ 5.3.3 MNIST Image Classification on an MLP
We next apply our new scheme to the MNIST digit classification dataset [28], a standard benchmark of computer vision applications. It consists of 60,000 training and 10,000 test examples of handwritten digits, stored as $28\times 28$ pixel grayscale images. As is customary, we normalize the data to mean 0.5 and standard deviation 0.5. We use a multi-layer perceptron (MLP) with two hidden layers with 800 and 300 nodes, respectively. While conventional neural network training is typically done with small batches, we use large batch sizes in our sampling experiments for now to prevent additional gradient noise from obscuring the effect of the (already noisy) stepsize adaptation. (For a preliminary study of the effect of varying batch size, refer to Appendix D.) Here we fix $B=10,000$. It is not our aim to demonstrate state-of-the-art performance for image classification of our new scheme. Rather, we want to illustrate the potential benefits of using adaptive-stepsize methods when exploring neural network loss landscapes compared to constant-stepsize methods. For the experiments in this and the next subsection, we use tranform kernel $\psi^{(1)}$ with $m=0.1$, $M=10$, $r=0.25$, and sample at $T=\gamma=1$. All trajectories are initialized through PyTorch’s default (the momenta being drawn from their invariant Gaussian measure). The monitor function is taken as $g(x,p)=N_{D}^{-1}\|\nabla U(x)\|^{2}$ with $N_{D}=60,000$ the number of examples in the training dataset. $\zeta$ is initialized to $g(0,0)$. Fig. 16 shows typical trajectories for BAOAB and SamAdams, where the former was run at the mean adaptive stepsize used by the latter. We observe a large loss spike in the constant-stepsize scheme which is not present for the adaptive scheme. Before and after the spike the dynamics seem to align to high degree, which implies that it is indeed the force-sensitive stepsize adaptation of SamAdams that prevents the spike from forming.
Figure 16: Single trajectory results for an MLP on MNIST. Top panel: training loss; middle panel: $\zeta$ dynamics; lower panel: adaptive stepsize $\Delta t$. BAOAB was run at the mean adaptive stepsize $\langle\Delta t\rangle$ of SamAdams. In this example we chose $\Delta\tau=0.0002$, and $\alpha=50$.
@@
While even the constant-stepsize scheme is able to recover from the instability, the appearance of spikes like this can lead to decreased performance in actual posterior sampling experiments, in which one averages over many different trajectories. To demonstrate this point, we draw 100 independent trajectories for both SamAdams and BAOAB, and average the resulting final accuracies. For SamAdams we use the same settings as in Fig. 16, i.e., $\Delta\tau=0.0002$ and $\alpha=50$. For BAOAB, we perform two experiments. In the first round, we set its learning rate to the mean of the stepsizes adopted by SamAdams, pooled from all 100 trajectories. In the second round, we take the mean of the pooled stepsizes again, but only consider the first 10% of $\Delta t$ samples adopted by SamAdams on each trajectory. The latter setup represents the idea that, for a fair comparison, the choice of stepsize for BAOAB should only be allowed to use information of the early stage of the SamAdams run. In contrast, using SamAdams’ mean adaptive stepsize uses information of all the stepsizes used by SamAdams, i.e., knowledge of the force evolution across the whole trajectory. This information would usually not be available to someone wanting to set the stepsize of a constant-stepsize scheme, so using the mean adaptive stepsize for the BAOAB runs gives the benefit of the doubt to BAOAB. At the same time, using the mean adaptive stepsize for BAOAB is the choice that leads to similar computational cost when run for the same number of iterations, which can also be the seen as the basis for a ’fair’ comparison. The final mean train and test accuracies together with 95%-confidence intervals are given by Table 1.666Note: these runs were performed without time averaging. We observe that SamAdams significantly outperforms both BAOAB setups.
ZBAOABZ BAOAB (mean of all $\Delta t$) BAOAB (mean of first 10% of $\Delta t$)
Train Accuracy (%) 94.0$\pm$0.1 92.3 $\pm$ 0.7 92.9 $\pm$ 0.5
Test Accuracy (%) 93.6$\pm$0.1 91.0 $\pm$ 0.7 91.5 $\pm$ 0.6
Table 1: Final accuracies averaged over 100 independent trajectories. Mean values and 95% confidence intervals. For the hyperparameters used and a description of the difference of the two BAOAB runs, see main text.

### v2

@@ 5.2.3 Beale Potential
Figure 10: Absolute errors of the means of four observables on the Beale potential against (mean) stepsize $\Delta t$. The ground truths for the coordinates were obtained via numerical quadrature. Simulation settings: $\gamma=1,\ T=3,\ \alpha=1,\ g(x,p)=0.1\|\nabla U\|^{2},\ \psi=\psi^{(1)}\text{with }m=0.1,\ M=10,\ r=0.25.$ See text for more information.
5.2.4 Auxiliary Variable Control Dynamics
Until now we have not yet explored the role of the hyperparameter $\alpha$ in (24) nor the influence of the scale of the monitor function $g(x,p)$. In the case of $g(x,p)=\Omega^{-1}\|\nabla U(x)\|^{s}$ (our choice for all of our experiments in this article), the scaling factor $\Omega$ governs how strongly $\zeta$ reacts to a given force value $\nabla U(x)$. Since the range of typical values of $\nabla U$ strongly depends on the problem at hand, the scaling parameter $\Omega$ is useful to adjust the range of $g(x,p)$ and hence the range of $\zeta$ and $\Delta t$. For the one- and two-dimensional examples considered thus far, setting $\Omega=1$ serves as a good starting point, whereas on the classification tasks of the the next section we use $\Omega=N_{\text{D}}$ with $N_{\text{D}}$ the size of the dataset. Since the force $\nabla U$ is written as a sum over the model likelihoods of all data points, this choice for $\Omega$ encourages $\zeta$-values close to the fixed point of the Sundman transform $\psi$. The hyperparameter $\alpha$ governs the damping of the $\zeta$-dynamics as well as its moving average behavior, i.e., how quickly past force values are ’forgotten’ (see Appendix C for more details).
As we have shown thus far, one fundamental benefit of the adaptive stepsize is the ability to use overall large stepsizes while using small stepsizes selectively. If the resulting $\langle\Delta t\rangle$ is larger than the maximum stable stepsize of a conventional scheme (assuming the accuracies remain acceptable), SamAdams leads to an improvement of stability and computational efficiency. An important question is then how difficult it is to tune the SamAdams-specific hyperparameters $\alpha$ and $\Omega$. To examine this, we chose the example of the star potential. We create a grid of different $(\alpha,\Omega)$-values and run SamAdams trajectories for each grid point. For each grid point, we experimentally find the stability threshold of SamAdams, i.e., the largest $\langle\Delta t\rangle$ that still leads to stable trajectories. We do this by running 100 independent trajectories for successively increasing values of $\Delta\tau$, the base stepsize, until the trajectories become unstable. We vary $\Delta\tau$ in $[0.04,0.08]$, a range in which the resulting $\langle\Delta t\rangle$-values become roughly comparable to the stability threshold of BAOAB for most ($\alpha,\Omega$)-combinations  [^4] BAOAB’s threshold was found to be $\Delta t=0.01275$, defined as the smallest stepsize for which 100 independent trajectories remain stable for $5\cdot 10^{6}$ iterations.. The number of iterations per tested $\Delta\tau$ is scaled via $N=N_{0}\Delta\tau_{0}/\Delta\tau$ with $N_{0}=2\cdot 10^{6}$ and $\Delta\tau_{0}=0.06$. If an $(\alpha,\Omega)$-combination leads to unstable trajectories even at the smallest $\Delta\tau$ tested, we deem the combination ‘highly unstable’. The results are shown in Fig. 11.
Figure 11: SamAdams stability thresholds $\langle\Delta t\rangle_{\text{max}}$ in dependency of $\alpha$ (attack rate) and $\Omega$ (scale coefficient of force norms) in the case of the star potential. Left: Plain $\langle\Delta t\rangle_{\text{max}}$. A well defined optimal zone appears around $\alpha=1$, $\Omega=100$. Right: Fraction of SamAdams threshold and BAOAB’s threshold $\Delta t^{\text{BAOAB}}_{\text{max}}$. The brown areas show parameter regions where BAOAB is more stable than SamAdams, the purple area regions where SamAdams is more stable. The red areas denote ’highly unstable’ regions (see main text). Other settings: $T=\gamma=1,\ s=2,\ \psi=\psi^{(1)}\text{ with }m=0.1,\ M=10,\ r=0.25$.
From the left-hand figure, we see how SamAdams’ stability threshold varies with $(\alpha,\Omega)$, with an optimum region visible for $\alpha\simeq 1$ and $\Omega^{-1}\simeq 0.01$. The highly-unstable area (red) is mainly due to too large values of $\alpha$. As shown in Appendix C, too large $\alpha$ will lead to $\Delta t=M\Delta\tau$. For our choice $M=10$ and the smallest $\Delta\tau$-value tested, $\Delta\tau=0.04$, this would correspond to a constant-stepsize scheme running at stepsize 0.4, more than 30 times larger than BAOAB’s stability threshold. Note that even for these highly unstable ($\alpha,\Omega)$-values, one could still enforce stability by picking $\Delta\tau$ and $M$ such that $M\Delta\tau$ is smaller than BAOAB’s threshold. On the right-hand side in Fig. 11, we plot the same grid but color according to the fraction of SamAdams’ stability threshold and BAOAB’s threshold. We observe that there is a large area in parameter space (shaded purple) in which SamAdams is more stable than BAOAB, with a remarkable improvement of up to 300%.
While these experiments show the need to pick admissible $(\alpha,\Omega)$-values, they also demonstrate that there is a wide range of settings in which SamAdams is more stable than BAOAB, reducing the need for time-consuming hyperparameter finetuning.
@@ 5.3 Higher-dimensional Models
We now explore several higher-dimensional cases using fixed stepsize Langevin as well as the SamAdams scheme. Our goal here is to show that there is a good basis for believing the method will be effective for applications in statistics and machine learning.
@@
Figure 12: Comparison of trajectory graphs in the $\theta,x_{1}$ projection obtained for different methods and stepsizes. Top: fixed stepsize BAOAB runs with $\Delta t$ increasing left to right until explosion is encountered at $\Delta t=0.1$. Bottom: results of different variable stepsize runs using SamAdams. While all these trajectories remain bounded, the last one shows some visible thickening in the $x$ direction, indicative of discretization bias.
We examined the stability of fixed stepsize (BAOAB) Langevin dynamics by running trajectories with a series of increasing timesteps $\Delta t=0.04,0.06,0.08,0.1$. The system was initialised with a large enough value of $\theta$ ($\theta_{0}=5$) and all $x_{i}=0$ so as to avoid any initial high gradient, and we sample at $\gamma=1$. The graphs of the $(\theta,x_{1})$-projection are shown in the upper row of Fig. 12. All simulations involved $N=10^{7}$ steps. We then ran SamAdams with various $\Delta\tau$, generating the figures shown in the second row, also using $10^{7}$ steps for each. The SamAdams monitor function is given by $g(x,p)\equiv\frac{1}{100}\|\nabla U(x)\|$, and the filter kernel is $\psi=\psi^{(1)}$ with $m=0.01,\ M=1.0,\ r=1$. We use $\alpha=1$ for the $\zeta$ dissipation and set $\zeta_{0}=0$. We label these runs by the mean stepsize used, which is reciprocally related to the computational work to reach a given fixed time. We observe that BAOAB only leads to accurate sampling at its lowest stepsize. Given the mild confining potential, the BAOAB excursions to large positive $x$ at stepsize greater than $0.04$ represent high energy (far from equilibrium) events; in a more complicated model those trajectories would lead to blow-up, or, perhaps more seriously, subtle degradation of cumulative averages; here the confining potential quickly returns them to the domain of interest. By contrast, SamAdams is stable and retains the trajectories within a similar size and shape region of the $(\theta,x_{1})$-plane, except for a slight increase in the width of the sampled region at the largest mean stepsize ($\langle\Delta t\rangle=0.1762$). This is noteworthy given the much larger stepsizes SamAdams is able to use. Since the largest stable mean stepsize of SamAdams is approximately four times the largest stable stepsize for Langevin dynamics with fixed stepsize, SamAdams promises substantial computational speedups.
@@
An important question is whether the time-series corresponding to different timesteps explore the space with similar efficiency. We settle this for the 9D funnel example in Fig. 14 where we see that relative to the elapsed time, the trajectories diffuse at similar rates. [^5] Computing autocorrelation functions and ESS requires first interpolating the non-uniformly spaced time-series data to a uniform mesh; this interpolation can be avoided when computing expectations by using the method of Section 2.1. Effective sample sizes per sample [18] are as shown in Table 1. These results indicate that the trajectories are similar in terms of the rates of exploration per unit time, but that the large mean stepsize SamAdams trajectories are much more efficient in terms of the samples gained per timestep taken.
Mean log posterior values (from 10M step runs) for the different stepsizes are also given in the table and suggest that bias may be becoming more noticeable at the larger stepsize. The result $-10.46$ for mean log posterior is accurate, verified using a small stepsize of $0.01$ and 1 billion steps. Accuracy (and ESS) appear to fall off at the largest fixed stepsize, with $\Delta t=0.02$ being approximately optimal. This can be compared to SamAdams with a mean stepsize of $\langle\Delta t\rangle=0.0998$, indicating an improvement of over 400% in sampling efficiency without any reduction in accuracy.
Note that with small stepsize $0.01$, fixed stepsize runs of 10M steps as in the table generate similar size errors (in this case due to Monte Carlo error due to the higher correlation of samples) to the largest variable stepsize integration (in that case, due to sampling bias), thus demonstrating a clear trade-off between sampling error and bias.
BAOAB SamAdams
@@
Table 1: Table showing the effective sample size per sample and the expected log posterior, for different variable stepsizes.
Figure 14: Autocorrelation functions with respect to time generated by different mean stepsizes. Left: autocorrelation of $\theta$; Right: autocorrelation of $x_{1}$. For each variable, the decay rates corresponding to different schemes are similar, regardless of the stepsize (and are similar for both the fixed and variable stepsize runs).
@@ 5.3.2 MNIST Image Classification on an MLP
We next apply our new scheme to the MNIST digit classification dataset [29], a standard benchmark of computer vision applications consisting of 60,000 training and 10,000 test examples of handwritten digits, stored as $28\times 28$ pixel grayscale images. The goal in this study is to illustrate the potential benefits of using adaptive-stepsize methods when exploring neural network loss landscapes compared to constant-stepsize methods. As is customary, we normalize the data to mean 0.5 and standard deviation 0.5. We use a multi-layer perceptron (MLP) with two hidden layers with 800 and 300 nodes, respectively. While conventional neural network training is typically done with small batches, we use large batch sizes ($B=10,000$) in the experiments reported here to prevent additional gradient noise from obscuring the effect of stepsize adaptation. (For study of the effect of varying batch size in logistic regression, refer to Appendix F.)
For the experiments in this and the next subsection, we use transform kernel $\psi^{(1)}$ with $m=0.1$, $\ M=10$, $\ r=0.25$. The monitor function is taken as $g(x,p)=N_{D}^{-1}\|\nabla U(x)\|^{2}$ with $N_{D}=60,000$ the number of examples in the training dataset. We use $\Delta\tau=0.0002$, $\alpha=50$, and sample at $T=\gamma=1$. All trajectories are initialized through PyTorch’s default (the momenta being drawn from their invariant Gaussian measure), with $\zeta_{0}=g(0,0)$.
Fig. 15 shows typical trajectories for BAOAB and SamAdams, where the former was run at the mean adaptive stepsize used by the latter. We observe a large loss spike in the constant-stepsize scheme which is not present for the adaptive scheme. Before and after the spike the dynamics seem to align to high degree, which implies that it is indeed the force-sensitive stepsize adaptation of SamAdams that prevents the spike from forming.
Figure 15: Single trajectory results for an MLP on MNIST. Top: Training loss for both BAOAB and SamAdams. The constant stepsize scheme BAOAB becomes unstable during training, observable as a large spike. Middle and bottom: Corresponding $\zeta$-dynamics and resulting adaptive stepsize $\Delta t$ of SamAdams.
@@
While even the constant-stepsize scheme is able to recover from the instability, the appearance of spikes like this can lead to decreased performance in actual posterior sampling experiments, in which one averages over many different trajectories. To demonstrate this point, we draw 100 independent trajectories for both SamAdams and BAOAB, and average the resulting final accuracies. For SamAdams we use the same settings as in Fig. 15. For BAOAB, we perform two experiments. In the first round, we set its learning rate to the mean of the stepsizes adopted by SamAdams, pooled from all 100 trajectories. In the second round, we take the mean of the pooled stepsizes again, but only consider the first 10% of $\Delta t$ samples adopted by SamAdams on each trajectory. The latter setup is based on the idea that, for a fair comparison, the choice of stepsize for BAOAB should only be allowed to rely on information of the early stage of the SamAdams run. In contrast, using SamAdams’ mean adaptive stepsize uses information of all the stepsizes used by SamAdams, i.e., knowledge of the force evolution across the whole trajectory. This information would usually not be available to someone wanting to set the stepsize of a constant-stepsize scheme, so using the mean adaptive stepsize for the BAOAB runs gives the benefit of the doubt to the fixed-stepsize method. At the same time, using the mean adaptive stepsize for BAOAB is the choice that leads to similar computational cost when run for the same number of iterations, which can also be seen as the basis for a ’fair’ comparison.
The final mean train and test accuracies together with 95%-confidence intervals are given by Table 2.666Note: these runs were performed without time-averaging. We observe that SamAdams significantly outperforms both BAOAB setups.
SamAdams BAOAB (mean of all $\Delta t$) BAOAB (mean of first 10% of $\Delta t$)
Train Accuracy (%) $\textbf{94.0}\pm\textbf{0.1}$ $92.3\pm 0.7$ $92.9\pm 0.5$
Test Accuracy (%) $\textbf{93.6}\pm\textbf{0.1}$ $91.0\pm 0.7$ $91.5\pm 0.6$
Table 2: Final accuracies averaged over 100 independent trajectories. Mean values and 95%-confidence intervals. For the hyperparameters used and a description of the difference of the two BAOAB runs, see main text.